# Multimodal ”Puppeteer”: An Exploration of Robot Teleoperation Via Virtual Counterpart with LLM-Driven Voice and Gesture Interaction in Augmented Reality

### v1

@@ Abstract
The integration of robotics and augmented reality (AR) holds transformative potential for advancing human-robot interaction (HRI), offering enhancements in usability, intuitiveness, accessibility, and collaborative task performance. This paper introduces and evaluates a novel multimodal AR-based robot puppeteer framework that enables intuitive teleoperation via virtual counterpart through large language model (LLM)-driven voice commands and hand gesture interactions. Utilizing the Meta Quest 3, users interact with a virtual counterpart robot in real-time, effectively “puppeteering” its physical counterpart within an AR environment. We conducted a within-subject user study with 42 participants performing robotic cube pick-and-place with pattern matching tasks under two conditions: gesture-only interaction and combined voice-and-gesture interaction. Both objective performance metrics and subjective user experience (UX) measures were assessed, including an extended comparative analysis between roboticists and non-roboticists. The results provide key insights into how multimodal input influences contextual task efficiency, usability, and user satisfaction in AR-based HRI. Our findings offer practical design implications for designing effective AR-enhanced HRI systems.
Index Terms:
@@
Robot puppeteer, teleoperation, virtual counterpart, multimodal interaction, user study, quantitative and qualitative analysis.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Robotics is rapidly reshaping numerous industries by facilitating automation, manufacturing, assistance, and collaboration across various domains [1, 2]. As intelligent robots become increasingly integrated into daily human-centered environments such as healthcare [3] and social interactions [4], it becomes essential to develop intuitive and natural interaction mechanisms. Traditionally, human-robot interaction (HRI) seeks to establish effective mutual communication between humans and robots to enhance task performance through collaboration or to improve user experience (UX) from a human-centered perspective [5, 6]. Augmented reality (AR)—a technology that overlays virtual elements onto the physical environment using specialized devices [7]—has emerged as a powerful and versatile tool to enhance HRI by providing real-time feedback [8, 9]. Historically, AR has been extensively employed for visualization due to its capability to directly present additional information in front of users. By bridging the physical and digital worlds, AR significantly improves usability, intuitiveness, situational awareness, and accessibility, making it increasingly valuable in the development of effective HRI systems today [10, 11, 12].
Within HRI systems, leveraging AR interfaces allows users to interact with robots in a more intuitive [13] and spatially aware manner [14]. This approach reduces cognitive load and operational difficulties [15, 16] associated with traditional interaction methods. Over recent years, numerous studies have explored the use of AR in robotic teleoperation, recognizing that additional virtual information can provide essential insights leading to more informative interactions and more robust robot control [17, 18]. For instance, AR has been successfully employed to improve robotic navigation efficiency [19] and enhance social interactions with dialogue-based robots [20], demonstrating the advantages offered by AR integration. Notably, most previous research on AR-assisted HRI or human-robot collaboration has predominantly focused on direct manipulation and control of physical robots. Among these contributions, the work by Haastregt et al. [21] stands out by introducing the first controller-based AR system for virtual robot teleoperation using a head-mounted display (HMD), enabling users to manipulate virtual robots via physical controllers.
An advancement in HRI has been the integration of multimodal interaction – combining various channels such as gestures, voice commands, haptic feedback, and spatial mapping—to facilitate enhanced communication between humans and robots [22, 23, 24]. However, the interdisciplinary field of multimodal interaction combined with AR-based virtual robotic teleoperation remains relatively unexplored. This approach is especially promising in extended reality (XR) environments, where users benefit from immersive virtual projections, improved spatial awareness, real-time feedback, and intuitive control methods when interacting with virtual robots, thus indirectly influencing physical robot systems. Compared to single-input interaction systems, multimodal interaction significantly narrows the gap between human intentions and robotic execution, leading to more seamless, efficient, and user-friendly interactions [25, 26, 27]. Previous research has demonstrated that integrating multiple modalities notably enhances overall UX and usability within HRI contexts [28, 29, 30]. Nonetheless, applying these multimodal approaches specifically to AR-based virtual robot control presents novel challenges, particularly concerning accurate voice command recognition, gesture detection, and synchronized robotic execution.
In this research, we present a multimodal robot “puppeteer” in AR – a novel framework that synchronizes physical robot and virtual counterpart for teleoperation, integrating large language model (LLM)-driven voice commands and hand gesture interactions within an AR environment. Unlike conventional approaches that treat virtual simulation and real-world robot control separately, our proposed system allows users to directly manipulate a virtual counterpart robot functioning as a graphical agent, translating and converting virtual commands seamlessly into the physical robotic movements. This puppeteer metaphor provides an intuitive, immersive, and precise robotic teleoperation experience, reducing the learning curve associated with complex robotic systems. By integrating voice and gesture-based interactions within an AR environment, our approach investigates the influence of interaction modalities on task performance, usability, and UX in the context of robotic puppeteering via virtual counterpart. We envision broader applications of our method in domains such as robot-assisted surgery, industrial automation, and operation in hazardous environments. The key contributions include:
@@ I Introduction
  • Develop an HMD AR-based puppeteer system supporting virtual-to-physical robot synchronized teleoperation with multimodal interaction: voice and hand gesture.
  • Evaluate the proposed system empirically, measuring user performance, usability, and UX.
  • Provide design implications with further insights between roboticists and non-roboticists.
II Related work
II-A XR for HRI
@@
Over the past decades, HRI has evolved significantly through the integration of XR—the umbrella term of virtual reality (VR), AR, and mixed reality (MR) [31]. These immersive technologies provide new paradigms for interaction beyond traditional screens and controllers. As early as 1999, Freund et al. [32] demonstrated that projective VR could support fundamental robotic functions such as task planning. In subsequent years, VR has been widely applied, for example, Matsas et al. [33] proposed a VR-based training system for human-robot collaboration in industrial settings. Similarly, Murnane et al. [34] utilized VR environments to collect interaction data for training verbal HRI models, while Villani et al. [35] employed VR as an evaluation tool to simulate complex HRI scenarios that would be difficult to reproduce in real-world environments. AR, which also encompasses MR, has gained broader attention in HRI due to its ability to overlay digital content onto the physical world, enabling real-time visualization, intuitive control, and spatial awareness. In collaborative HRI, Chacko et al. [36] developed a smartphone-based AR interface for pick-and-place tasks, allowing users to visually select target objects within a shared workspace. Michalos et al. [37] introduced an AR system for industrial co-working environments, providing assembly guidance, safety cues, and live production feedback, validated in an automotive use case. For teleoperation, Hedayati et al. [38] presented AR interfaces that improved visual feedback and user performance during aerial robot inspections. Walker et al. [39] further extended this by introducing a virtual surrogate drone in AR to enhance control efficiency, reduce cognitive load, and support multitasking—showing measurable improvements across both novice and expert user groups.
II-B Multimodal Interaction in XR for HRI
@@
Multimodal interaction – leveraging a combination of input modalities such as gesture, voice, gaze, and haptics – has become a promising area in XR-enabled HRI for enhancing control precision, accessibility, and user engagement. In early VR-based research, Savage et al. [40] developed a VR system that enables users to simulate controlling a mobile robot in virtual environments. The robot responds to voice and simple gesture commands, with contextual reasoning enhancing command recognition. More recently, the focus has shifted to AR-based multimodal systems. Park et al. [41] proposed a hands-free HRI method using head gestures, eye gaze, and deep learning to enable efficient robot control via indirect object manipulation and a digital twin system. Chan et al. [42] developed an AR platform integrating speech, gestures, gaze, electromyography, and tactile feedback for robot trajectory programming, emphasizing low-attention, force-aware interaction. Szczurek et al. [43] designed a multimodal AR interface for collaborative intervention in hazardous environments, using head-mounted displays to support shared workspaces and reliable control.
Figure 1: The proposed multimodal robot puppeteer system with AR.
II-C Robot Puppeteer and Research Gap
Despite advancements in AR-based HRI and multimodal interaction, research on virtual-to-physical robot control using a puppeteering metaphor in AR remains limited. Barakonyi et al. [20] explored animated agents in AR to facilitate seamless control across physical and virtual domains, while Walker et al. [39] demonstrated the use of virtual drone surrogates to improve teleoperation and reduce user stress. More recently, van Haastregt et al. [21] introduced the first AR-based puppeteer system that allows users to control a physical robot for simple movements via manipulation of a virtual counterpart using a physical HMD controllers in hands. However, several gaps remain unaddressed: (1) the integration of multimodal interaction – specifically combining gesture and voice – in the context of AR-based virtual-to-physical robot puppeteering has yet to be explored; (2) the evaluation of contextual usability and UX within such systems, especially among roboticists. Our work aims to address these gaps by presenting an interactive, multimodal AR system that enables users to interact with a virtual counterpart robot using a natural, puppeteer-like metaphor, thereby manipulating a corresponding physical robot in real-time. Unlike a digital twin, our virtual robot functions as a standalone graphical proxy in AR, without real-time bi-directional data flow synchronization with a physical robot – similar to the setup described in [21].
@@ II-C Robot Puppeteer and Research Gap
III Research design
III-A The Robot Puppeteer System
The developed AR-based robotic puppeteer system centers a virtual counterpart robot, which mirrors a physical robot in real-time with Meta Quest 3. Through continuous synchronization, the physical robot replicates the movements of its virtual counterpart. As illustrated in Fig. 1, the system is built upon the Franka robotic arm – a widely used 7-degree-of-freedom collaborative robot designed for manipulation tasks, with a 3D-printed yellow gripper mounted at the end-effector  [^1] https://franka.de/. The virtual counterpart robot is rendered in light green within the AR environment and projected directly in the user’s field of view to facilitate intuitive interaction. The experiment subjects used in this study include a few pre-defined zones within the joint range of the physical robot based on six colored cubes (see Section  IV-C for details). The connection between the physical robot and the virtual robot (the AR environment) was achieved via Unity-Robot Operating System (ROS) TCP  [^2] https://github.com/Unity-Technologies/ROS-TCP-Endpoint. The technical realization of the puppeteer system is displayed in Figure  2. The position and orientation of the virtual robot are defined by user configuration in the real study, as described in the following sections.
III-B The Voice and Gesture Interaction
@@
III-C Apparatus
IV User study
IV-A Overview of the Study
We designed and conducted a within-subject comparative user study to evaluate the effectiveness of our robot puppeteering system. Participants were asked to complete two identical cube pattern-matching tasks under two distinct conditions: gesture only (GO) and voice + gesture (VG) (Fig.4 and Fig.4). During the task, participants were instructed to pick up six colored cubes (red, yellow, blue, orange, purple, and green) from a designated starting zone, grasp them, and place them in the target zone to match a predefined color cell (Fig. 4). The cube pick-and-place task is a well-established benchmark in robotic teleoperation research, widely recognized for its effectiveness in assessing interaction methods [44, 45, 46]. Additionally, pattern-matching tasks have been commonly used in HCI and psychology studies to measure task accuracy and cognitive load [47, 48, 49]. In the GO condition, users relied exclusively on hand gestures for task execution, while in the VG condition, they could use a combination of voice commands and gestures. The independent variable in this study was the interaction modality (GO vs. VG). We initially hypothesized:
@@ IV-A Overview of the Study
(c) The experiment setup with pre-defined zones and colored cubes.
@@ IV-C Procedure
Participants began by completing a pre-study questionnaire designed to collect background information, including prior experience with VR/AR technologies and robotics. They then proceeded to the task execution phase, which involved a cube pick-and-place and pattern-matching task to be completed within a 5min (300s) time limit. An example of the user’s AR view is shown in Fig. 4, where the instruction interface appears floating on the right side, the microphone on the left, and the virtual robot is positioned directly in front of the user after spawning, which remained still throughout the whole study. Participants were instructed to maintain simultaneous visual awareness of both the virtual and physical robots throughout the task. Although virtual reference cubes were displayed in the AR environment, participants were informed that these elements were intended solely for contextual reference and not for interaction. During the task, participants interacted with the virtual counterpart robot for control and manipulation, while visually focusing on the physical robot to perform the pick-and-place actions in the puppeteer metaphor. The task began by moving the robot from its initial location to the starting zone, where colored cubes were placed in correspondingly colored cells (Fig.4). Using the yellow gripper mounted on the robot’s end-effector (Fig.4), participants grasped each cube and placed it into the matching cell in the target zone to complete a color pattern.
@@
Table I: Self-developed measurement metrics in the user study.
Metric Category Definition/Explanation Range
Performance Metrics
@@
Perceived Intent to Us ”I would like to use this system for robot control.” [1, 7]

### v2

@@ Abstract
The integration of robotics and augmented reality (AR) offers promising opportunities to enhance human–robot interaction (HRI) by making teleoperation more transparent, spatially grounded, and intuitive. We present a head-mounted AR ”puppeteer” framework in which users control a physical robot via interacting with its virtual counterpart robot using large language model (LLM)-driven voice commands and hand-gesture interaction on the Meta Quest 3. In a within-subject user study with 42 participants performing an AR-based robotic pick-and-place pattern-matching task, we compare two interaction conditions: gesture-only (GO) and combined voice+gesture (VG). Our results show that GO currently provides more reliable and efficient control for this time-critical task, while VG introduces additional flexibility but also latency and recognition issues that can increase workload. We further explore how prior robotics experience shapes participants’ perceptions of each modality. Based on these findings, we distill a set of evidence-based design guidelines for AR puppeteer metaphoric robot teleoperation, implicating multimodality as an adaptive strategy that must balance efficiency, robustness, and user expertise rather than assuming that additional modalities are universally beneficial. Our work contributes empirical insights into how multimodal (voice+gesture) interaction influences task efficiency, usability, and user experience in AR-based HRI.
keywords:
@@
Robot puppeteer, teleoperation, virtual counterpart, multimodal interaction, user study, quantitative and qualitative analysis
1 Introduction
@@
Figure 1: The AR-based multimodal robot “puppeteer” system used in this study, enabling voice and gesture interaction for teleoperation.
Robotics is rapidly reshaping numerous industries by facilitating automation, manufacturing, assistance, and collaboration across various domains [ribeiro2021robotic, hentout2019human, xia2024shaping]. As intelligent robots become increasingly integrated into daily human-centered environments such as healthcare [olaronke2017state, kyrarini2021survey] and social interactions [richert2016socializing, zhang2024human], it becomes essential to develop intuitive and natural interaction mechanisms. Traditionally, human-robot interaction (HRI) seeks to establish effective mutual communication between humans and robots to enhance task performance through collaboration or to improve user experience (UX) from a human-centered perspective [hopko2021effect, apraiz2023evaluation, zhang2025mind]. Augmented reality (AR)—a technology that overlays virtual elements onto the physical environment using specialized devices [zhou2008trends, zhang2021novel]—has emerged as a powerful and versatile tool to enhance HRI by providing real-time feedback [saeedi2018navigating, suzuki2022augmented]. Historically, AR has been extensively employed for visualization due to its capability to directly present additional information in front of users [zhang2022initial, nowak2021augmented, zhang2021supporting]. By bridging the physical and digital worlds, AR significantly improves usability, intuitiveness, situational awareness, and accessibility, making it increasingly valuable in the development of effective HRI systems today [yoo2025study, bambusek2019combining, gaschler2014intuitive].
Within HRI systems, leveraging AR interfaces allows users to interact with robots in a more intuitive [chong2009robot, zhang2023playing] and spatially aware manner [lunding2023ar]. This approach reduces cognitive load and operational difficulties [walker2018communicating, moya2023augmented] associated with traditional interaction methods. Over recent years, numerous studies have explored the use of AR in robotic teleoperation, recognizing that additional virtual information can provide essential insights leading to more informative interactions and more robust robot control [arevalo2021assisting, makhataeva2020augmented]. For instance, AR has been successfully employed to improve robotic navigation efficiency [mohareri2011autonomous] and enhance social interactions with dialogue-based robots [barakonyi2004agents], demonstrating the advantages offered by AR integration. Notably, most previous research on AR-assisted HRI or human-robot collaboration has predominantly focused on direct manipulation and control of physical robots. Among these contributions, the work by Haastregt et al. [van2024puppeteer] stands out by introducing the first controller-based AR system for virtual robot teleoperation using a head-mounted display (HMD), enabling users to manipulate virtual robots via physical controllers.
An advancement in HRI has been the integration of multimodal interaction – combining various channels especially gestures, voice commands, and haptic feedback – to facilitate enhanced communication between humans and robots [salem2011friendly, torta2015evaluation, okita2011multimodal]. However, the interdisciplinary field of multimodal interaction combined with AR-based virtual robotic teleoperation remains relatively unexplored. This approach is especially promising in extended reality (XR) environments, where users benefit from immersive virtual projections, improved spatial awareness, real-time feedback, and intuitive control methods when interacting with virtual robots, thus indirectly influencing physical robot systems. Compared to single-input interaction systems, multimodal interaction significantly narrows the gap between human intentions and robotic execution, leading to more seamless, efficient, and user-friendly interactions [ali2019design, oviatt1999ten, lee2013usability, zhang2023see]. Previous research has demonstrated that integrating multiple modalities notably enhances overall UX and usability within HRI contexts [lunghi2019multimodal, lazzeri2014development, cao2023investigating]. Nonetheless, applying these multimodal approaches specifically to AR-based virtual robot control presents novel challenges, particularly concerning accurate voice command recognition, gesture detection, and synchronized robotic execution [zhang2024vision].
In this research, we investigate how multimodal interaction—specifically voice and hand gesture—shapes teleoperation of a physical robot via its virtual counterpart in head-mounted AR, with a focus on expertise-aware target groups. We present an AR “puppeteer” prototype that synchronizes a physical robot and its virtual twin for teleoperation, integrating large language model (LLM)-driven voice commands with hand-gesture interaction. Rather than treating virtual simulation and real-world robot control as separate workflows, our system lets users directly manipulate the virtual counterpart as a graphical control agent, translating these virtual manipulations into physical robot motion. By involving an empirical study, we tend to examine how different modalities affect task performance, usability, and user experience, and how these effects vary with users’ prior expertise. Based on observed findings, we derive actionable, expertise-sensitive design implications for multimodal AR teleoperation, and discuss the implications for the HCI/HRI community. The key contributions include:
  • Present an HMD AR puppeteer teleoperation prototype that instantiates gesture-only (GO) and voice+gesture (VG) conditions for virtual-to-physical robot interaction.
  • Report a within-subject user study (N=42) comparing GO and VG for a structured pick-and-place teleoperation task, analyzing task performance, usability, and UX.
  • Derive empirical, expertise-aware design implications for AR puppeteer-style voice+gesture teleoperation, clarifying the use of voice and gesture based on participant feedback, and propose multimodal interaction insights.
2 Related work
2.1 XR for HRI
@@
Over the past decades, HRI has evolved significantly through the integration of XR—the umbrella term of virtual reality (VR), AR, and mixed reality (MR) [williams2018virtual]. These immersive technologies provide new paradigms for interaction beyond traditional screens and controllers. As early as 1999, Freund et al. [freund1999projective] demonstrated that projective VR could support fundamental robotic functions such as task planning. In subsequent years, VR has been widely applied, for example, Matsas et al. [matsas2017design] proposed a VR-based training system for human-robot collaboration in industrial settings. Similarly, Murnane et al. [murnane2021simulator] utilized VR environments to collect interaction data for training verbal HRI models, while Villani et al. [villani2018use] employed VR as an evaluation tool to simulate complex HRI scenarios that would be difficult to reproduce in real-world environments. AR, which also encompasses MR, has gained broader attention in HRI due to its ability to overlay digital content onto the physical world, enabling real-time visualization, intuitive control, and spatial awareness. In collaborative HRI, Chacko et al. [chacko2019augmented] developed a smartphone-based AR interface for pick-and-place tasks, allowing users to visually select target objects within a shared workspace. Michalos et al. [michalos2016augmented] introduced an AR system for industrial co-working environments, providing assembly guidance, safety cues, and live production feedback, validated in an automotive use case. For teleoperation, Hedayati et al. [hedayati2018improving] presented AR interfaces that improved visual feedback and user performance during aerial robot inspections. Walker et al. [walker2019robot] further extended this by introducing a virtual surrogate drone in AR to enhance control efficiency, reduce cognitive load, and support multitasking—showing measurable improvements across both novice and expert user groups.
2.2 Multimodal Interaction in XR for HRI
@@
Multimodal interaction – leveraging a combination of input modalities such as gesture, voice, gaze, and haptics – has become a promising area in XR-enabled HRI for enhancing control precision, accessibility, and user engagement. In early VR-based research, Savage et al. [savage1998virbot] developed a VR system that enables users to simulate controlling a mobile robot in virtual environments. The robot responds to voice and simple gesture commands, with contextual reasoning enhancing command recognition. More recently, the focus has shifted to AR-based multimodal systems. Krupke et al. [krupke2018comparison] developed an MR system that lets users control co-located virtual industrial robot arms for pick-and-place with comparing heading+speech versus pointing+speech for pick selection. Park et al. [park2021hands] proposed a hands-free HRI method using head gestures, eye gaze, and deep learning to enable efficient robot control via indirect object manipulation and a digital twin system. Chan et al. [chan2022multimodal] developed an AR platform integrating speech, gestures, gaze, electromyography, and tactile feedback for robot trajectory programming, emphasizing low-attention, force-aware interaction. Szczurek et al. [szczurek2023multimodal] designed a multimodal AR interface for collaborative intervention in hazardous environments, using head-mounted displays to support shared workspaces and reliable control.
2.3 Robot Puppeteer and Research Gap
Regarding controlling robots using AR featured virtual robots, Duan et al. [duan2023ar2] introduced an iOS pipeline: AR2-D2 to turn phone videos of a person manipulating objects into robot demonstrations via an AR Franka overlay and motion planning, but removing the need for a real robot during data collection and targeting behavior-cloning pipelines. Similarly, Wang et al. [wang2024eve] extended the work by proposing an iOD AR app – EVE – to let users set and edit AR waypoints on iOS to gather training trajectories and verify them, emphasizing data quality and downstream policy results, not live control of any physical robots. Despite these advancements, research on virtual-to-physical robot control using a puppeteering metaphor in AR remains limited. Barakonyi et al. [barakonyi2004agents] explored animated agents in AR to facilitate seamless control across physical and virtual domains, while Walker et al. [walker2019robot] demonstrated the use of virtual drone surrogates to improve teleoperation and reduce user stress. Krupke et al. [krupke2018comparison] proposed an HMD-based AR concept in which a visualized robotic arm is co-located with the real robot, allowing users to control the virtual arm to rehearse tasks before executing them on the physical robot. In terms of virtual-to-physical synchronization, van Haastregt et al. [van2024puppeteer] introduced the first AR-based puppeteer system that allows users to control a physical robot for simple movements via manipulation of a virtual counterpart using a physical HMD controllers in hands, while Zhang et al. [zhang2025llm] upgraded similar system by introducing the controller-free interaction with simple LLM integration for voice command.
However, several gaps remain. First, the use of multimodal interaction, especially comparing voice and gesture interaction, in AR-based virtual-to-physical robot puppeteering has received limited attention. Second, there is a lack of empirical evaluations that jointly examine task performance, contextual usability, and UX in such systems. Third, prior work offers little design-oriented guidance that accounts for differences between experts/non experts. Our work addresses these gaps by investigating how multimodality (voice and gesture) shapes AR-based robot teleoperation in a puppeteer-like metaphor. Unlike a full digital twin, our virtual robot in AR functions as a standalone graphical proxy – without live bidirectional data flow with the physical robot – similar to the setups described in [van2024puppeteer, zhang2025llm], but extended with an expertise-aware, multimodal evaluation.
3 Research design
3.1 The Multimodal Robot Puppeteer System
The developed AR-based robotic puppeteer system centers a virtual counterpart robot, which mirrors a physical robot in real-time with Meta Quest 3. Through continuous synchronization, the physical robot replicates the movements of its virtual counterpart. As illustrated in Fig. 1, the system is built upon the Franka robotic arm – a widely used 7-degree-of-freedom collaborative robot designed for manipulation tasks, with a 3D-printed yellow gripper mounted at the end-effector  [^1] https://franka.de/. The virtual counterpart robot is rendered in light green within the AR environment and projected directly in the user’s field of view to facilitate intuitive interaction. The experiment subjects used in this study include a few pre-defined zones within the joint range of the physical robot based on six colored cubes (see Section 4.3 for details). The connection between the physical robot and the virtual robot was achieved via Unity-Robot Operating System (ROS) TCP  [^2] https://github.com/Unity-Technologies/ROS-TCP-Endpoint. The technical realization of the puppeteer system is displayed in Figure 2. The position and orientation of the virtual robot are defined by user configuration in the real study, as described in the following sections.
@@ 3.1 The Multimodal Robot Puppeteer System
3.2 The Voice and Gesture Interaction
@@ 3.3 Apparatus
4 User study
4.1 Overview of the Study
We conducted a within-subject comparative user study to evaluate the effectiveness of our robot puppeteering system. Participants were asked to complete two identical cube pattern-matching tasks where they were instructed to pick up six colored cubes (red, yellow, blue, orange, purple, and green) from a designated starting zone, grasp them, and place them in the target zone to match a predefined color cell (Fig. 4(c)). The cube pick-and-place task is a well-established benchmark in robotic teleoperation research, widely recognized for its effectiveness in assessing interaction methods [fang2014novel, ostanin2020human, pandey2014towards]. Additionally, pattern-matching tasks have been commonly used in HCI and psychology studies to measure task accuracy and cognitive load [seaborn2010exploring, swan2015matching, nickerson1973visual]. To better match the puppeteer nature of our system and the hands-on demands of pick-and-place manipulation, we implemented two interaction conditions: gesture-only (GO) and voice + gesture (VG) (Fig.4(a) and Fig.4(b)). A voice-only condition was not included in this paper, as it offers limited spatial precision and was deemed impractical for this task. In the GO condition, users relied exclusively on hand gestures for task execution, while in the VG condition, they could use a combination of voice commands and gestures. The independent variable in this study was the interaction modality (GO vs. VG). We initially hypothesized:
@@ 4.1 Overview of the Study
(c) The setup with pre-defined zones and colored cubes.
@@ 4.3 Procedure
Participants began by completing a pre-study questionnaire designed to collect background information, including prior experience with VR/AR technologies and robotics. They then proceeded to the task execution phase, which involved a cube pick-and-place and pattern-matching task to be completed within a 5min (300s) time limit. An example of the user’s AR view is shown in Fig. 4(b), where the instruction interface appears floating on the right side, the microphone on the left, and the virtual robot is positioned directly in front of the user after spawning, which remained still throughout the whole study. Participants were instructed to maintain simultaneous visual awareness of both the virtual and physical robots throughout the task. Although virtual reference cubes were displayed in the AR environment, participants were informed that these elements were intended solely for contextual reference and not for interaction. During the task, participants interacted with the virtual counterpart robot for control and manipulation, while visually focusing on the physical robot to perform the pick-and-place actions in the puppeteer metaphor. The task began by moving the robot from its initial location to the starting zone, where colored cubes were placed in correspondingly colored cells (Fig.4(c)). Using the yellow gripper mounted on the robot’s end-effector (Fig.4(d)), participants grasped each cube and placed it into the matching cell in the target zone to complete a color pattern.
@@
Table 1: Self-developed measurement metrics in the user study.
[…]
@@
Perceived Intent to Use ”I would like to use this system for robot control.” [1, 7]